### v2

# Toward Any-to-Any Emotion Voice Conversion using Disentangled Diffusion Framework

Abstract
Emotional Voice Conversion (EVC) aims to modify the emotional expression of speech for various applications, such as human-machine interaction. Previous deep learning-based approaches using generative adversarial networks and autoencoder models have shown promise but suffer from quality degradation and limited emotion control. To address these issues, a novel diffusion-based EVC framework with disentangled loss and expressive guidance is proposed. Our method separates speaker and emotional features to maintain speech quality while enhancing emotional expressiveness. Tested on real-world and acted-out datasets, the approach achieved significant improvements in emotion classification accuracy for both in-the-wild and act-out datasets and showed reduced distortion compared to state-of-the-art models.
Index Terms:
Emotion Voice Conversion, Diffusion, Disentanglement, Guidance
I Introduction
Emotional Voice Conversion (EVC), which focuses on artificially modifying the emotional expression of speech signals, offers potential applications such as enhancing the naturalness of human-machine interaction [1, 2]. By increasing the controllability over complex factors, including voice quality, emotional expressiveness, speaker traits, and linguistic contents, for human speech generation, we can gradually perform one-to-one, many-to-many, and finally, any-to-any emotion voice conversion, which is applicable to convert the emotion of unseen speakers. With the effectiveness of deep learning techniques, building and evaluating the EVC model on specific target speakers or small-scaled parallel emotional speech has gained much success in previous works [3, 4, 5, 6].
When addressing various EVC issues, most of the deep learning methods models lie in the two realms, which are generative adversarial networks [3, 4] and autoencoder-based models [5, 6]. GAN-based methods leverage adversarial mechanisms to learn direct mappings between data distributions of different emotional classes. Autoencoder-based approaches tackle this issue by introducing a disentanglement mechanism that separates linguistic and speaker identity features from emotional representations, enabling better control over emotion conversion. However, these works are suboptimal in converting the source emotion to the target emotion and can distort the converted voice, i.e., degrading the generated voice’s naturalness and emotion quality.
The diffusion model has recently garnered significant attention for its generative capabilities of high-quality samples in multiple application scenarios [7, 8, 9]. A most recent emotion conversion work, EmoConv-Diff [10] demonstrated a statistically significant improvement in intensity controllability compared to their previous method [11] while achieving great quality. However, the controllability of valence is still challenging, and it is unclear whether the speaker and emotion information are learned and extracted from the corresponding entangled representation. To tackle the problems, we design our framework with disentanglement in two separate aspects to generate better emotional characteristics while sustaining high speech quality comparable to original source utterances. First, we incorporate the disentangle loss into the diffusion EVC model training process. Second, with the reversed diffusion process, we further design the expressive guidance mechanism to enhance the expressiveness of the target emotion and mitigate the distortion of the source speaker trait.
In this work, different from working in the controlled scenarios such as [3, 4, 5, 6], we aim to solve any-to-any emotional voice conversion by proposing a novel diffusion framework design with disentangled loss and expressive guidance. To demonstrate the effectiveness, we train our model on un-parallel real-world speech emotion data (MSP-Podcast v1.10 [12]) and rigorously evaluate our method on two different corpora, including the real-world scenario (MSP-Podcast v1.11) and act-out scenario (ESD [1]), with the unseen environment, speakers and linguistic contents. Both objective and subjective evaluations show that our proposed method effectively converts emotion in speech for in-the-wild data with 21% ECA improvement on MSP and 32% ECA improvement on ESD. Moreover, with objective metrics of 3.718 in UTMOS, 3.600 in SIG, and subjective metrics of 4.024 in nMOS, and 3.159 in sMOS, our experiment results demonstrate that our proposed any-to-any EVC can generate emotional speech comparable to state-of-the-art EVC with significantly less distortion.
Figure 1: Overview of the proposed diffusion-based any-to-any emotion voice conversion framework.
II Methodology
II-A Proposed Method
This work aims to solve the problem of any-to-any emotion voice conversion. Given a pair of $X_{src}:=g(c_{src},spk_{src},emo_{src})$ and reference $X_{ref}:=g(c_{ref},spk_{ref},emo_{ref})$ speech utterances, where each utterance are composed with linguistic content $c$, speaker identity $spk$ and emotion information $emo$ and, $g(.)$ is a generative process, our proposed method $G$ aims to perform the conversion process $\hat{X}=G(c_{src},spk_{src},emo_{ref})$ that preserve both content and speaker identity while transforming emotion from $emo_{src}$ to $emo_{ref}$. The any-to-any setting requires the source and reference speech utterance to be completely unseen in the training set.
Figure 1 shows the overall framework of our proposed method, which comprises a set of encoders modeling each component, a diffusion-based decoder, a disentanglement mechanism, and a guidance method. We use pre-trained HiFiGAN vocoder [13] to convert the mel spectrogram generated by the proposed framework back to the time domain signal.
II-A1 Encoders
Three pre-trained encoders are used to capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
Phoneme Encoding: To encode linguistic content $\bar{X}$, we adapt pre-trained transformer-based encoder from [14] to convert input mel-spectrograms $X_{0}$ into speaker and emotion independent “average-voice” mel features that replaced each phoneme-level mel feature with the corresponding average phoneme-level mel features.
Speaker Encoding: To encode the speaker identity $Z_{s}\in\mathbb{R}^{256}$, we use a pretrained speaker verification model [15] adapted from [14].
Emotion Encoding: To encode emotional information $Z_{e}\in R^{1024}$, we use a SSL-based SER system adapted from [16] that was built by fine-tuning the Wav2Vec2-Large-Robutst [17] network on the MSP-Podcast (v1.7) dataset [12].
To disentangle speaker and emotion representations, we encode the corresponding disentangled representations as $\hat{Z_{s}}=h_{s}(Z_{s})$ and $\hat{Z_{e}}=h_{e}(Z_{e})$, where $h_{s}$ and $h_{e}$ are linear transformations with learnable parameters.
II-A2 Diffusion Decoder
We employ the diffusion framework based on stochastic differential equations (SDE) from [14], conditioned on given representations $\bar{X},Z_{s},Z_{e}$ to generate high-quality speech. The diffusion process gradually transforms the real sample ${X_{0}}$ into ${X_{t}}$ with time-step $t\in[0,1]$ that terminates at average-voice mel-spectrogram $\bar{X}$ when $t=1$ by adding Gaussian noise in a forward process; and generates ${X_{0}}$ from $\bar{X}$ by removing the corresponding score estimation $s_{\theta}(X_{t},t,\bar{X},\hat{Z_{s}},\hat{Z_{e}})$ in a reverse process. The $s_{\theta}$ with parameter $\theta$ is trained by minimizing mean square error loss $\mathcal{L}_{diff}$ between added noise and $s_{\theta}$.
II-A3 Expressive Guidance
To amplify the effectiveness of the diffusion model on the converted speech, we further design the expressive guidance method that aims to manage the reversed diffusion process with positive and negative direction scores. During the inference stage, we modified $s_{\theta}$ with $\hat{s}_{\theta}$ as follows:
$\hat{s}_{\theta}$ $=s_{\theta,neg}+\lambda_{EG}(s_{\theta,pos}-s_{\theta,neg})$ (1)
$\lambda_{EG}$ with the value >1 controls the intensity of this guidance method and pushes the generation process away from the negative condition but toward the positive condition. For any-to-any emotion voice conversion, the positive condition takes the source linguistic content $c_{src}$, the source speaker identity $spk_{src}$, and the reference emotion information $emo_{ref}$; On the other hand, the negative condition can be either changing $spk_{src}$ to $spk_{ref}$ for $EG_{spk}$, $emo_{ref}$ to $emo_{src}$ for $EG_{emo}$ or both for $EG_{spk,emo}$, where $EG$ stands for the proposed expressive guidance method.
II-A4 Disentangled Loss
To reduce the correlation between different speech representations, specifically emotion information and speaker identity, we minimize MI loss between representations $\mathcal{L}_{MI}=\hat{I}(\hat{z}_{s},\hat{z}_{e})$, where $\hat{I}$ represents the unbiased estimation for vCLUB as described in [18]. The MI loss has shown to be effective in disentangling between different speech representations in several studies [19, 20].
To further preserve speaker identity and emotion information residing in the representations after disentanglement, we use two auxiliary supervised models that 1) predict speaker identity from disentangled speaker representation $\hat{z}_{s}$, and 2) predict emotion label (Neutral, Angry, Happy, and Sad) and emotion attributes (Arousal and Valence) from disentangled emotion representation $\hat{z}_{e}$. These models are trained to minimize loss $\mathcal{L}_{style}$ where the negative log-likelihood loss is used for the categorical prediction task and the concordance correlation coefficient loss is used for the regression task.
In addition to $L_{diff}$ for training diffusion-based decoder, we follow [10] to use a mel-spectrogram recontruction loss $\mathcal{L}_{rec}$ that measures the $\mathcal{L}_{1-norm}$ between $X_{0}$ and $\hat{X}_{0}$, where $\hat{X}_{0}$ is the single-step approximation relying on $X_{t},\bar{X},s_{\theta}$ using Tweedie’s formula [21]. We use $\lambda_{rec}=(1-t^{2})$ adapt from [10] to reduce the importance of the loss with $X_{t}$ containing more Gaussian noise for larger t. The final objective function for our proposed method is as follows
$\mathcal{L}_{Total}=\mathcal{L}_{diff}+\lambda_{MI}\mathcal{L}_{MI}+\lambda_{style}\mathcal{L}_{style}+\lambda_{rec}\mathcal{L}_{rec}$ (2)
where $\lambda_{MI}$ and $\lambda_{guide}$ are hyparameters to control the importance of respective loss.
TABLE I: Objective evaluation of different training and inference schemes of proposed methods for any to any voice conversion. We also report the percentage of improvement of ECA compared to the baseline method
Methods MSP-Podcast (1.11) ESD
UTMOS SECS DNSMOS ECA UTMOS SECS DNSMOS ECA
SIG OVRL SIG OVRL
Source 2.613 1.000 3.365 2.902 0.627 3.919 1.000 3.463 3.178 0.960
Baseline ($\mathcal{L}_{diff}$) 2.362 0.821 3.511 3.075 0.701 3.712 0.794 3.550 3.289 0.493
[…]
III Experimental Setup and Results
III-A Experimental Setup
III-A1 Implementation Details
Our proposed methodology is trained on in-the-wild MSP-Podcast corpus (v1.10) [12] that contains real podcast recordings (16kHz, 1ch) with emotional expressions segmented in utterances. We select 53685 utterances labeled with four emotion and emotion attributes from 1385 labeled speakers. We adopted pre-trained model parameters from [14] and fine-tuned it on MSP-Podcast for 368k iterations with a batch size of 32. The Adam optimizer with a learning rate of $1\times 10^{-4}$ is used to update the trainable model parameters. We set $\lambda_{MI}=0.1$ and $\lambda_{style}=1$ during training, and set $\lambda_{EG}=1.25$ for expressive guidance during inference.
III-A2 Evaluation Setup
We evaluate our methods on both in-the-wild datasets, MSP-Podcast (v1.11), with real-world scenarios and act-out dataset, ESD [22] with high-quality recordings. We randomly sample 100 utterances of each emotion category with unseen speakers from both datasets to conduct the following experiments.
First, we perform any-to-any emotion voice conversion that includes all of the transformations between angry, happy, sad, and neutral, besides transforming from emotional speech to neutral one. The experiments are conducted on both in-the-wild datasets (MSP-Podcast) and high-quality act-out datasets (ESD). We compared methods under different training schemes, i.e., using only $\mathcal{L}_{diff}$ or using $\mathcal{L}_{Total}$ in equation 2. We then apply the proposed expressive guidance method on the model trained with $\mathcal{L}_{Total}$. We compared $EG_{spk}$, $EG_{emo}$, and $EG_{spk,emo}$ with different settings of negative condition that replace the representation of positive condition corresponding to either speaker identity $spk$, emotion information $emo$ or both. Second, We compared our proposed method with $EG_{spk,emo}$ against five different models, i.e. CycleGAN-EVC [3], StarGAN-EVC [4], Seq2Seq-EVC [5], Emovox [6] and Prosody2Vec [23] following conversion samples based on ESD presented in Prosody2Vec [^1] https://leyuanqu.github.io/Prosody2Vec/. Unlike ours, models in comparison are trained on act-out ESD datasets, while only Prosody2Vec utilized both predominant and in-the-wild datasets. The audio samples are available on our demo page [^2] https://henrychou36.github.io/Any-to-Any-EVC/.
III-A3 Evaluation Metric
For both experiments, we incorporate non-intrusive objective evaluation, i.e., UTMOS [24] for natureliness, DNSMOS [25] for speech quality (SIG) and overall signal quality (OVRL). Both methods are designed to predict the mean opinion score (MOS) of subjective listening tests. To access speaker similarity, speaker embedding cosine similarity (SECS) between extracted embeddings of source and generated speech based on Resemblyzer [15] is used. For emotion classification accuracy (ECA), we utilized a speech emotion recognition (SER) model pre-trained on both MSP-Podcast and ESD based on emotion embedding from [16]. For the second experiment, in addition to objective evaluation, we conducted a subjective assessment with 14 subjects evaluating 63 converted or target utterances using a 5-point scale ranging from 1 to 5 to assess speech quality, naturalness, and emotion similarity between synthesis speech and the target utterances. We report the mean opinion scores with a 95% confidence interval for speech quality (MOS), naturalness (nMOS), and emotion similarity (sMOS). The subjects are also required to label the primary emotion for ECA. The evaluation of the first and second experiments are presented in Table I and Table II, separately.
TABLE II: Objective and subjective evaluation of our proposed method and comparison models under unseen test dataset.
Method framework Objective Subjective
UTMOS SECS DNSMOS ECA MOS nMOS sMOS ECA
SIG OVRL
Target 3.606 0.816 3.429 3.155 1.000 4.484$\pm$0.120 4.397$\pm$0.149 4.976$\pm$0.027 0.794
CycleGan-EVC one-to-one 2.687 0.839 3.296 2.990 0.444 3.659$\pm$0.201 3.635$\pm$0.190 2.476$\pm$0.246 0.270
StarGan-EVC many-to-many 3.128 0.884 3.461 3.190 0.222 4.024$\pm$0.178 4.119$\pm$0.182 2.325$\pm$0.236 0.206
Seq2Seq-EVC many-to-many 1.903 0.663 3.301 2.957 0.444 2.595$\pm$0.217 2.032$\pm$0.175 2.008$\pm$0.212 0.103
Emovox many-to-mnay 2.381 0.698 3.234 2.930 0.333 2.683$\pm$0.202 2.191$\pm$0.158 2.508$\pm$0.235 0.325
Prosody2Vec many-to-many 2.482 0.730 3.071 2.717 0.889 3.095$\pm$0.201 2.778$\pm$0.180 3.484$\pm$0.229 0.603
Our Method any-to-any 3.718 0.757 3.600 3.341 0.889 4.024$\pm$0.171 4.318$\pm$0.149 3.159$\pm$0.258 0.508
III-B Experimental Results
From the result presented in Table I and Table II, the proposed method with both disengagement mechanism and expressive guidance method of $EG_{spk,emo}$ significantly improves in ECA with at least 21% compared to the underlying baseline model, and achieves high naturalness (3.718 UTMOS and 4.024 MOS) and high quality (3.600 SIG and 4.318 nMOS) samples that are target-comparable while having effective emotion conversion in terms of emotion similarity (3.159 sMOS) compared to other baseline models. This demonstrates that proposed any-to-any frameworks can generate emotional speech with less distortion.
Evaluating the disentanglement mechanism with Table I, we find out that the disentanglement framework itself significantly improves at least 16% ECA for both datasets compared to the baseline model. Comparing guidance methods under different settings of unwanted representation with only disentanglement mechanism, we find out that by utilizing unwanted source emotion information $e_{src}$, the generated speech is more emotional with at least 25% ECA increments, while damages naturalness, speaker similarity, and speech quality. On the other hand, using unwanted reference speaker identity $s_{ref}$ decreases the emotion accuracy while improving other criteria slightly. The downside of both methods can be alleviated by jointly considering unwanted speaker and emotion information, resulting in a 21% ECA improvement for MSP-Podcast and 32% ECA improvement for ESD. This overall result shows that the guidance method can control the expressiveness of emotion voice conversion in either speaker identity or emotion information.
Since previous works, such as EmoConv-Diff, primarily focus on evaluating the effectiveness of diffusion models in controlling emotion intensity, we aim to assess both emotion intensity and valence based on the confusion matrix in Figure 2. For the baseline diffusion model, while it can generate angry and sad speech distinct from each other in terms of emotion intensity, it struggles to generate differentiable speech that varies in valence but has similar intensity—such as angry and happy speech. This indicates that, although the underlying diffusion model can control emotion intensity similarly to EmoConv-Diff, controlling valence remains challenging. However, by incorporating our disentanglement mechanism and expressive guidance method, our model generates more emotionally distinct speech that can be better differentiated from non-target emotions.
Comparing source speech and synthesis speech regardless of different training and inference schemes, the synthesis results have overall better speech and audio quality, which shows that noise can be alleviated through speech decomposition and reconstruction. However, the naturalness is restricted by the naturalness of source utterances.
As the preliminary work to address any-to-any EVC problem, we achieve comparable or even better performance than one-to-one and many-to-many frameworks in terms of naturalness, quality, and expressiveness based on the result presented in Table II. Compared to the GAN-based methods having the highest speaker similarity with unrecognizable emotion and Prosody2Vec which generates the most emotional speech while suffering from low naturalness and audio quality, our method generates relatively distortionless emotional speech with utterances from unseen speakers.
Figure 2: Confusion Matrix comparison of different methods for MSP-Podcast. X-axis: classified labels. Y-axis: desired labels.
IV Conclusion and future work
It is important to develop ”any-to-any” EVC for real-world applications. In this work, we propose an any-to-any emotion voice conversion that combines a disentanglement mechanism and expressive guidance and provides thorough evaluation with both objective and subjective tests over both in-the-wild and act-out datasets. We show that our proposed framework can effectively convert speech into different emotions while having high speech and audio quality. Moreover, compared to other EVC frameworks in the control environment, our method generates distortionless emotional speech. We also address the challenge of controllability over valence, which will be further evaluated and developed for our future works.

### v3

@@
 To address these issues, a novel diffusion-based EVC framework with disentangled loss and expressive guidance is proposed.
-Our method separates speaker and emotional features to maintain speech quality while enhancing emotional expressiveness.
+Our method separates speaker and emotional features to maintain speech quality while improving emotional expressiveness.
 Tested on real-world and acted-out datasets, the approach achieved significant improvements in emotion classification accuracy for both in-the-wild and act-out datasets and showed reduced distortion compared to state-of-the-art models.
@@
 I Introduction
-Emotional Voice Conversion (EVC), which focuses on artificially modifying the emotional expression of speech signals, offers potential applications such as enhancing the naturalness of human-machine interaction [1, 2].
+Emotional Voice Conversion (EVC), which focuses on artificially modifying the emotional expression of speech signals, offers potential applications such as improving the naturalness of human-machine interaction [1, 2].
 By increasing the controllability over complex factors, including voice quality, emotional expressiveness, speaker traits, and linguistic contents, for human speech generation, we can gradually perform one-to-one, many-to-many, and finally, any-to-any emotion voice conversion, which is applicable to convert the emotion of unseen speakers.
-With the effectiveness of deep learning techniques, building and evaluating the EVC model on specific target speakers or small-scaled parallel emotional speech has gained much success in previous works [3, 4, 5, 6].
-When addressing various EVC issues, most of the deep learning methods models lie in the two realms, which are generative adversarial networks [3, 4] and autoencoder-based models [5, 6].
+With the effectiveness of deep learning techniques, building and evaluating the EVC model with specific target speakers or small-scale parallel emotional speech has gained much success in previous works [3, 4, 5, 6].
+When addressing various EVC issues, most deep learning methods models lie in the two realms, which are generative adversarial networks [3, 4] and autoencoder-based models [5, 6].
 GAN-based methods leverage adversarial mechanisms to learn direct mappings between data distributions of different emotional classes.
@@
 However, these works are suboptimal in converting the source emotion to the target emotion and can distort the converted voice, i.e., degrading the generated voice’s naturalness and emotion quality.
-The diffusion model has recently garnered significant attention for its generative capabilities of high-quality samples in multiple application scenarios [7, 8, 9].
+The diffusion model has recently gained significant attention for its generative capabilities of high-quality samples in multiple application scenarios [7, 8, 9].
 A most recent emotion conversion work, EmoConv-Diff [10] demonstrated a statistically significant improvement in intensity controllability compared to their previous method [11] while achieving great quality.
 However, the controllability of valence is still challenging, and it is unclear whether the speaker and emotion information are learned and extracted from the corresponding entangled representation.
-To tackle the problems, we design our framework with disentanglement in two separate aspects to generate better emotional characteristics while sustaining high speech quality comparable to original source utterances.
+To tackle the problems, we design our framework with disentanglement in two separate aspects to generate better emotional characteristics while sustaining high speech quality comparable to source utterances.
 First, we incorporate the disentangle loss into the diffusion EVC model training process.
 Second, with the reversed diffusion process, we further design the expressive guidance mechanism to enhance the expressiveness of the target emotion and mitigate the distortion of the source speaker trait.
-In this work, different from working in the controlled scenarios such as [3, 4, 5, 6], we aim to solve any-to-any emotional voice conversion by proposing a novel diffusion framework design with disentangled loss and expressive guidance.
-To demonstrate the effectiveness, we train our model on un-parallel real-world speech emotion data (MSP-Podcast v1.10 [12]) and rigorously evaluate our method on two different corpora, including the real-world scenario (MSP-Podcast v1.11) and act-out scenario (ESD [1]), with the unseen environment, speakers and linguistic contents.
-Both objective and subjective evaluations show that our proposed method effectively converts emotion in speech for in-the-wild data with 21% ECA improvement on MSP and 32% ECA improvement on ESD.
-Moreover, with objective metrics of 3.718 in UTMOS, 3.600 in SIG, and subjective metrics of 4.024 in nMOS, and 3.159 in sMOS, our experiment results demonstrate that our proposed any-to-any EVC can generate emotional speech comparable to state-of-the-art EVC with significantly less distortion.
+In this work, different from working in controlled scenarios such as [3, 4, 5, 6], we aim to solve any-to-any emotional voice conversion by proposing a novel diffusion framework design with disentangled loss and expressive guidance.
+To demonstrate the effectiveness, we train our model on nonparallel real-world speech emotion data (MSP-Podcast v1.10 [12]) and rigorously evaluate our method on two different corpora, including the real-world scenario (MSP-Podcast v1.11) and act-out scenario (ESD [13]), with the unseen environment, speakers and linguistic contents.
+Both objective and subjective evaluations show that our proposed method effectively converts emotion in speech for in-the-wild data with 21% ECA improvement on MSP and 32% ECA improvement on ESD datasets.
+Furthermore, with objective metrics of 3.718 in UTMOS, 3.600 in SIG, and subjective metrics of 4.024 in nMOS, and 3.159 in sMOS, our experimental results demonstrate that our proposed any-to-any EVC can generate emotional speech comparable to state-of-the-art EVC with significantly less distortion.
 Figure 1: Overview of the proposed diffusion-based any-to-any emotion voice conversion framework.
@@
 Figure 1 shows the overall framework of our proposed method, which comprises a set of encoders modeling each component, a diffusion-based decoder, a disentanglement mechanism, and a guidance method.
-We use pre-trained HiFiGAN vocoder [13] to convert the mel spectrogram generated by the proposed framework back to the time domain signal.
+We utilize pre-trained HiFiGAN vocoder [14] to convert the mel spectrogram generated by the proposed framework back to the time domain signal.
 II-A1 Encoders
-Three pre-trained encoders are used to capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
+Three pre-trained encoders capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
 Phoneme Encoding: To encode linguistic content $\bar{X}$, we adapt pre-trained transformer-based encoder from [15] to convert input mel-spectrograms $X_{0}$ into speaker and emotion independent “average-voice” mel features that replaced each phoneme-level mel feature with the corresponding average phoneme-level mel features.
@@
 $\mathcal{L}_{Total}=\mathcal{L}_{diff}+\lambda_{MI}\mathcal{L}_{MI}+\lambda_{style}\mathcal{L}_{style}+\lambda_{rec}\mathcal{L}_{rec}$ (2)
-where $\lambda_{MI}$ and $\lambda_{guide}$ are hyparameters to control the importance of respective loss.
+where $\lambda_{MI}$ and $\lambda_{style}$ are hyparameters to control the importance of respective loss.
 TABLE I: Objective evaluation of different training and inference schemes of proposed methods for any to any voice conversion.
@@
 We randomly sample 100 utterances of each emotion category with unseen speakers from both datasets to conduct the following experiments.
-First, we perform any-to-any emotion voice conversion that includes all of the transformations between angry, happy, sad, and neutral, besides transforming from emotional speech to neutral one.
+First, we performed any-to-any emotion voice conversion that includes all of the transformations between angry, happy, sad, and neutral, besides transforming from emotional speech to neutral one.
 The experiments are conducted on both in-the-wild datasets (MSP-Podcast) and high-quality act-out datasets (ESD).
@@
 III-A3 Evaluation Metric
-For both experiments, we incorporate non-intrusive objective evaluation, i.e., UTMOS [24] for natureliness, DNSMOS [25] for speech quality (SIG) and overall signal quality (OVRL).
+For both experiments, we incorporate non-intrusive objective evaluation, i.e., UTMOS [24] for naturelness, DNSMOS [25] for speech quality (SIG) and overall signal quality (OVRL).
 Both methods are designed to predict the mean opinion score (MOS) of subjective listening tests.
@@
 Seq2Seq-EVC many-to-many 1.903 0.663 3.301 2.957 0.444 2.595$\pm$0.217 2.032$\pm$0.175 2.008$\pm$0.212 0.103
-Emovox many-to-mnay 2.381 0.698 3.234 2.930 0.333 2.683$\pm$0.202 2.191$\pm$0.158 2.508$\pm$0.235 0.325
+Emovox many-to-many 2.381 0.698 3.234 2.930 0.333 2.683$\pm$0.202 2.191$\pm$0.158 2.508$\pm$0.235 0.325
 Prosody2Vec many-to-many 2.482 0.730 3.071 2.717 0.889 3.095$\pm$0.201 2.778$\pm$0.180 3.484$\pm$0.229 0.603
@@
 We show that our proposed framework can effectively convert speech into different emotions while having high speech and audio quality.
-Moreover, compared to other EVC frameworks in the control environment, our method generates distortionless emotional speech.
+Moreover, our method generates distortionless emotional speech compared to other EVC frameworks in the control environment.
 We also address the challenge of controllability over valence, which will be further evaluated and developed for our future works.